OpenStudy (anonymous):

Which of the following expressions is equal to 5 - 2m? 7 - 2(m + 1) 7 - (2m - 5) 7 - 2(m - 2)

OpenStudy (usukidoll):

we can distribute the negative... and distribute the -2 until we find an equation that equals 5-2m

OpenStudy (anonymous):

You would open up the brackets for each of the choices. Like your previous problem.

OpenStudy (anonymous):

im just copying and pasting.....

OpenStudy (anonymous):

okay so its not a

OpenStudy (anonymous):

Did you work it out? Distribute the -2

OpenStudy (anonymous):

i did i got 7 - 2

OpenStudy (usukidoll):

wait a second ... distribute the -2 for 7-2(m+1)

OpenStudy (usukidoll):

what's -2 x m what's -2 x 1

OpenStudy (anonymous):

What happen to the m?

OpenStudy (anonymous):

i keep forget that the letter is a 1

OpenStudy (anonymous):

-2

OpenStudy (oleg3321):

so it would be the 3rd option

OpenStudy (oleg3321):

do you understand how to do it now? @audieeeuniverse

OpenStudy (usukidoll):

what's -2 x m what's -2 x 1 = -2 @audieeeuniverse

OpenStudy (anonymous):

-2

OpenStudy (usukidoll):

you're forgetting the m

OpenStudy (anonymous):

-2m?

OpenStudy (oleg3321):

yes so it would be the 3rd option and you are forgeting the 2m

OpenStudy (usukidoll):

yes so now we have 7-2m-2

OpenStudy (oleg3321):

yes -2m

OpenStudy (usukidoll):

now what's 7-2 ?

OpenStudy (anonymous):

5

OpenStudy (oleg3321):

idk what is it lol 5

OpenStudy (usukidoll):

yes it is 5.. so we have 5-2m... isn't that the same as the question from earlier?

OpenStudy (oleg3321):

@UsukiDoll give me a medal please

OpenStudy (anonymous):

Practice makes perfect. Don't forget your variables!

OpenStudy (usukidoll):

it should be 7-2(m+1) = 5-2m 7-2m-2=5-2m 5-2m=5-2m

OpenStudy (usukidoll):

no @oleg3321 you gave the wrong answer... the third option is already out thanks to -2 x -2 which is 4

OpenStudy (oleg3321):

oh my mad. and ok it ok i messed up a little

OpenStudy (anonymous):

sooo its 7 - (2m - 5)

OpenStudy (usukidoll):

k I'll give a medal anyway since you've admitted to your errors @oleg3321

OpenStudy (oleg3321):

thank you i appretiate it

OpenStudy (anonymous):

7 - 2 (1 +1) 5 times 2 7

OpenStudy (usukidoll):

that one shouldn't be there .. and 5 x 2 = 10 here's what I wrote again 7-2(m+1) = 5-2m 7-2m-2=5-2m 5-2m=5-2m

OpenStudy (anonymous):

no i got this i get it because 7-2 is 5 then 1+19because of the m and then that makes it 7 -2 ( m + 1)